OpenStudy (anonymous):

The base of a triangle is 6 cm greater than the height. The area is 20 cm^2. Find the height and length of the base

OpenStudy (anonymous):

a = 1/2b*h 20=(x +6)(x)/2 20 = (x^2+6x)/2 40=x^2+6x 40 = (x+3)^2 -9 49=(x+3)^2 7=x+3 x=4 b= 4+6, h = 4. a=1/2b*h, 10*4/2 = 20.

OpenStudy (anonymous):

Or base = 10

OpenStudy (anonymous):

height = 4